According to indexmundi.com China produces the largest amount of ammonia with 45,520 in thousand metric ton. The profitability of the ammonia production is around 90% of the total ammonia production cost which is about Php 10,000 per ton.
Refer to the balanced chemical equation: N2(g) + 3H2(g) ➡️ 2NH3(g)

How much nitrogen and hydrogen were used in the production of ammonia in China?
1
Expert's answer
2020-11-25T02:18:39-0500

m(H2)=m(NH3)*3Mr(H2)/2Mr(NH3)=45520*3*2/(2*17)=8,033 thousand metric ton;

m(N2)=m(NH3) - m(H2)=45,520-8,033=37,487 thousand metric ton;